OpenStudy (anonymous):

Use elimination to solve the system. What is the value of x? 4x + 7y = -14 and 8x + 5y = 8

jimthompson5910 (jim_thompson5910):

Hint: Multiply everything in the first equation by -2 to go from 4x + 7y = -14 to -8x - 14y = 28

jimthompson5910 (jim_thompson5910):

So you now have the two equations -8x - 14y = 28 8x + 5y = 8

OpenStudy (anonymous):

That's solving for x first and then plugging it back to get y. I thought that the OP wanted to do it from that one step.

jimthompson5910 (jim_thompson5910):

you can go a number of ways, you can multiply the top equation by 5 and the bottom equation by -7, then add

OpenStudy (anonymous):

im just getting more confused

OpenStudy (anonymous):

I believe the answer is -14.

OpenStudy (anonymous):

these are the options. 3.5 -1.5 8 -4

OpenStudy (anonymous):

Oh then i bet the answer is -5

jimthompson5910 (jim_thompson5910):

Do you see how I went from 4x + 7y = -14 to -8x - 14y = 28

jimthompson5910 (jim_thompson5910):

it's not -5

jimthompson5910 (jim_thompson5910):

please don't give away the (incorrect) answer

OpenStudy (anonymous):

sorry, if you would just give her the right answer, i would be able to go on with my day. As would she

jimthompson5910 (jim_thompson5910):

I'm providing hints and trying to walk her through it (without giving it all away)

OpenStudy (anonymous):

No matter how many hints your provide she doesn't care. she just wants the answer

jimthompson5910 (jim_thompson5910):

well that's not what this site is for, I'm sry

jimthompson5910 (jim_thompson5910):

if you keep this behavior up jefferyflow123go you will be reported (and possibly suspended/banned)

OpenStudy (anonymous):

its hard for me too understand..im trying. i learn better in a class room thats why i am having trouble with my online class

jimthompson5910 (jim_thompson5910):

its ok, you don't have to understand everything at once katie2212

OpenStudy (anonymous):

but i have to understand if i want to finishe this exam.

OpenStudy (anonymous):

Sorry for hekping. My mom always taught me to hekp and have pride in my work, so trying to be of assistance. I didn;t realize that you can be the only genious here. If you were that smart you'd know that she doesn't understand your hints and just wants the answer.

jimthompson5910 (jim_thompson5910):

that's not helping though, you may think it is...but it's not

jimthompson5910 (jim_thompson5910):

if you want, you can provide the steps so she can learn from them

OpenStudy (anonymous):

okay will you guys please stop arguing and just help me!

jimthompson5910 (jim_thompson5910):

ok I'll give steps on this one and hopefully it will be enough to answer similar ones, sound good?

OpenStudy (anonymous):

IS your refrigderator running ? IF SO GO CATCH IT

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

My momma once told me, a fast runna' is the best runnna'

OpenStudy (anonymous):

jeffery if youre not going to help please get out.

OpenStudy (anonymous):

BUT IM JUST TRYNA HELP YOU. ILL CATCH YOUR FRIDGE FR YOU.

OpenStudy (anonymous):

What did the bison say to his son when he left for college? Bi-son

jimthompson5910 (jim_thompson5910):

4x + 7y = -14 ... Start with the first equation -2(4x + 7y) = -2*(-14) ... Multiply both sides by -2 -8x - 14y = 28 ... Distribute and multiply --------------------- So 4x + 7y = -14 is the same as -8x - 14y = 28 --------------------- This means that the system 4x + 7y = -14 8x + 5y = 8 becomes -8x - 14y = 28 8x + 5y = 8 Now add the two equations: -8x - 14y = 28 8x + 5y = 8 -------------- 0x - 9y = 36 which means that -9y = 36 Solve for y to get -9y = 36 y = 36/(-9) y = -4 Now that we know that y = -4, we can use this to find x. Pick on any equation, plug in y = -4 and solve for x 4x + 7y = -14 4x + 7(-4) = -14 4x - 28 = -14 4x = -14 + 28 4x = 14 x = 14/4 x = 7/2 x = 3.5 So the value of x is x = 3.5

OpenStudy (anonymous):

IM SO PUN-NY

OpenStudy (anonymous):

IM A MaTH WIZ

OpenStudy (anonymous):

Thank you Jim

OpenStudy (anonymous):

yanno wizard like Harry Potter

jimthompson5910 (jim_thompson5910):

The idea here is to get the x coefficients equal and opposite. Then you add the equations. Doing so will cancel out the x terms so you can solve for y

jimthompson5910 (jim_thompson5910):

then you use that value of y to find x

jimthompson5910 (jim_thompson5910):

you've been reported jefferyflow123go, have a nice day

OpenStudy (anonymous):

you too good sir(:

jimthompson5910 (jim_thompson5910):

good luck with that one, I don't spam and harass others so your claim will be without merit